OpenStudy (anonymous):

Find the pct to the nearest thousandth. Games won- 15; Games played- 25

OpenStudy (anonymous):

Multiply 25*4 = 100 Multiply 15*4 = 60 So 60% of the games were won. Hope this helps

OpenStudy (anonymous):

Where did you get the 4?

OpenStudy (anonymous):

Percent out of 100

OpenStudy (anonymous):

25*4 = 100% 15*4 = 60 Therefore 60/100

OpenStudy (anonymous):